OpenStudy (anonymous):

multiply square root 7 * square root 3

OpenStudy (anonymous):

In decimal form, it would be 4.58

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b}=\sqrt{ab}\] assuming a>=0 and b>=0

OpenStudy (mathmagician):

\[\sqrt{7}*\sqrt{3}=\sqrt{21}\]

OpenStudy (anonymous):

I was going to say what @Dockworker and @mathmagician said, but they beat me to the punch, so I'll explain why \(\sqrt a\sqrt b=\sqrt{ab}\mid\forall\left(a,b\geq0\right)\). That's because the plane of complex numebrs is undefined for such an operation, so that you can run into paradoxes, the most obvious example being: \(\sqrt{-1}\sqrt{-1}=i^2=-1\) versus \(\sqrt{-1}\sqrt{-1}=\sqrt{1}=1\).